A random sample of 900 members is found to have a mean of 4.45 CM can it be reasonably regarded as a sample from a large population whose mean is 5 cm and variance is 4 cm square obtain the 95% confidence limit for the mean in the population

1
Expert's answer
2022-03-18T12:43:11-0400

"95\%CI=(\bar x-t_{0.025,n-1}\frac{\sigma}{\sqrt{n}}, \bar x+t_{0.025,n-1}\frac{\sigma}{\sqrt{n}})="

"=(4.45-1.963\frac{2}{\sqrt{900}}, 4.45+1.963\frac{2}{\sqrt{900}})="

"=(4.319, 4.581)."